OpenStudy (anonymous):

Gasoline is pouring into a cylindrical tank of radius 4 feet. When the depth of the gasoline is 6 feet, the depth is increasing at 0.3 ft/sec. How fast is the volume of gasoline changing at that instant? Round your answer to 3 decimal places! The volume is changing at a rate of _______ feet^3 per second. @ganeshie8 do you get this one? :/

ganeshie8 (ganeshie8):

start by writing the Volume of cylinder formula..

ganeshie8 (ganeshie8):

\(V = \pi r^2 h = \pi \times 4^2\times h \\ ~~~= 16\pi h\)

OpenStudy (anonymous):

umm is it this? :/ πr^2h = v ?

OpenStudy (anonymous):

ohh okay haha :P

ganeshie8 (ganeshie8):

yes :) differentiate both sides with.respect.to \(t\): \(V = 16\pi h\) \(\dfrac{dV}{dt} = ?\)

OpenStudy (anonymous):

ummm 16pi ?

ganeshie8 (ganeshie8):

Note that we're NOT differentiating with.respect.to "h",

ganeshie8 (ganeshie8):

we're differentiating with.respect.to "t"

ganeshie8 (ganeshie8):

so you need to use chain rule

ganeshie8 (ganeshie8):

\(V = 16\pi h\) \(\dfrac{dV}{dt} = 16 \pi \dfrac{dh}{dt}\)

OpenStudy (anonymous):

ahh not sure how to do that here :( so what happens after that? :/

ganeshie8 (ganeshie8):

And you're given `the depth is increasing at 0.3 ft/sec. `

ganeshie8 (ganeshie8):

that means \(\dfrac{dh}{dt} = 0.3\)

ganeshie8 (ganeshie8):

plug that in and simplify

OpenStudy (anonymous):

ahh okay so 16pi(0.3) = 15.07964474 ? so rounded to three decimal places, we get 15.101 ? did i round that correctly ? :/ or should it be 15.081? :/

ganeshie8 (ganeshie8):

http://prntscr.com/3cuij0

OpenStudy (anonymous):

whoah awesome! thank you!! :D

ganeshie8 (ganeshie8):

np :)